### v1

# A Comprehensive Framework for Semantic Similarity Detection Using Transformer Architectures and Enhanced Ensemble Techniques

Abstract
Detecting AI-generated text, especially in short-context documents, is difficult because there is not enough context for accurate classification. This paper presents a new teacher-student model that uses domain adaptation and data augmentation to solve these problems. The teacher model, which combines DeBERTa-v3-large and Mamba-790m, learns semantic knowledge through domain-specific fine-tuning. The student model handles short-context text more efficiently. The system uses a Mean Squared Error (MSE) loss function to guide the student’s learning, improving both accuracy and efficiency. Also, data augmentation methods like spelling correction and error injection make the model more robust. Experimental results show that this approach works better than baseline methods, proving its usefulness for real-time AI-generated text detection and other text classification tasks.
Index Terms:
AI-generated text detection, teacher-student model, data augmentation, short-context document classification, DeBERTa-v3-large
I Introduction
The rise of AI-generated content, driven by language models like GPT-3 and BERT, has created problems for content moderation and text classification. Detecting machine-generated text is important for many uses, including combating misinformation and verifying academic work. A main problem is detecting AI-generated text in short-context documents, where traditional models often fail due to limited context. Short-context texts, such as tweets, headlines, or excerpts, lack the longer structures that models usually rely on for classification. As a result, even small changes in the text can have a big impact on performance.
To solve this problem, we propose a teacher-student model that combines the strengths of large pre-trained models and short-context text classification. The teacher model, made up of DeBERTa-v3-large and Mamba-790m, is fine-tuned with domain adaptation. This helps it capture more detailed semantic features and better detect AI-generated content. The teacher model helps guide the student model, which is optimized for short-context handling and efficiency. The student model is smaller and more efficient, trained to predict the teacher’s output using a Mean Squared Error (MSE) loss function. This allows it to benefit from the teacher’s knowledge while being fast enough for real-time use.
We also improve the model’s robustness with data augmentation. By applying text changes like spelling correction, character removal, and adding errors, the model learns to handle real-world variations. These augmentations help the model generalize better, making it more accurate in detecting AI-generated text, even in noisy or imperfect conditions. The teacher-student framework, domain adaptation, and data augmentation together create a model that performs well in short-context text classification, where other models struggle.
II Related Work
Detecting AI-generated text has become an important research area because of the growth of large language models (LLMs) and generative adversarial networks (GANs). Yan et al.[1] discuss generative LLMs, focusing on the challenges they present for distinguishing AI-generated text from human-written text. These challenges are also present with GANs, which Gui et al.[2] discuss in terms of their applications for content generation. GANs are particularly used for creating fake text, a problem addressed by Zellers et al.[3] in their study on defending against fake news generated by neural networks.
To improve detection of AI-generated text, Chakraborty et al.[4] review different methods for identifying machine-generated content, looking at how various AI models perform. A major challenge is the domain specificity of text, as models trained on general datasets may not work well in specialized areas. For example, SciBERT, a model for scientific text, has been successful in detecting AI-generated academic papers by recognizing domain-specific language patterns[5]. Dehaerne et al.[6] also explore machine learning for detecting machine-generated code, highlighting difficulties in identifying such content.
Some studies focus on user interactions with AI-generated content. Lu[7] suggests using decision trees and TF-IDF to improve chatbot user satisfaction, which can also help detect AI-generated dialogue. Li[8] looks at how multimodal data can improve product recommendations, a method that could also be used to detect AI-generated content by combining different data sources.
Text summarization models, like those used by Liu and Lapata[9], have shown promise in detecting AI-generated text by analyzing the structure and coherence of the content. Schick and Schütze[10] study few-shot learning, which could be used to detect subtle linguistic patterns of AI-generated text.
The problem of detecting AI-generated text involves many areas, such as fake news, code, and academic papers. By using domain-specific models, machine learning techniques, and better detection algorithms, progress can be made in identifying AI-generated content.
III Methodology
Semantic similarity between phrases is a critical task in the patent search and examination process. This paper presents a comprehensive study on advanced deep learning models and ensemble strategies for semantic similarity detection. By integrating transformer-based architectures, bidirectional LSTM layers, and novel tricks like adversarial weight perturbation and linear attention pooling, we achieve state-of-the-art performance. Additionally, we introduce dynamic target grouping and fine-tuned ensemble methods to boost diversity and robustness, ensuring superior generalization. The pipline of model is shown in Fig 1.
Figure 1: The pipline of transformer-based architectures model.
III-A Transformer Backbone
We utilize DeBERTa-v3-large as the primary feature extractor:
$X_{bert}=\text{DeBERTa}(X_{input}),$ (1)
where $X_{input}$ represents the tokenized input sequence. The DeBERTa backbone provides contextualized embeddings, leveraging disentangled attention to capture fine-grained relationships. Additionally, freezing the embedding layers during fine-tuning stabilizes training and reduces overfitting, as the semantic similarity task involves short text sequences.
III-B Bidirectional LSTM Enhancement
To capture sequential dependencies and enrich feature representation, a Bi-LSTM layer is appended to the transformer outputs:
$X_{lstm}=\text{Bi-LSTM}(X_{bert}),$ (2)
where $X_{lstm}$ combines forward and backward dependencies. Adversarial Weight Perturbation (AWP) is introduced during the second epoch to enhance robustness by simulating adversarial scenarios, ensuring that the Bi-LSTM learns more generalizable features. The pipline of LSTM Enhancement is shown in Fig 2.
Figure 2: The pipline of Bidirectional LSTM Enhancement.
III-C Linear Attention Pooling
For dimensionality reduction and improved focus on key features, linear attention pooling is applied:
$X_{pool}=\sum_{t=1}^{T}\alpha_{t}\cdot X_{lstm,t},$ (3)
where $\alpha_{t}$ are learned attention weights, and $T$ is the sequence length. Dynamic target shuffling during each training step augments this module by exposing the pooling layer to diverse target sequences, enhancing generalization. The linear attention pooling is shown in Fig 3.
Figure 3: The linear attention pooling.
III-D Fully Connected Layer
The final representation is passed through a fully connected layer to compute the similarity score:
$Y_{pred}=\text{FC}(X_{pool}).$ (4)
Differentiated learning rates are applied, with a lower learning rate ($2e^{-5}$) for the transformer and a higher rate ($1e^{-3}$) for the LSTM and fully connected layers. This strategy ensures efficient optimization while preserving the pre-trained knowledge of the transformer.
III-E Alternative Model Architectures
To enhance ensemble diversity, we incorporated additional architectures, each tailored to leverage specific strengths:
III-E1 Electra-Based Models
Electra models, pre-trained with a replaced token detection (RTD) objective, complement the transformer backbone by capturing finer-grained semantic nuances. The model is formulated as:
$X_{electra}=\text{Electra}(X_{input}),$ (5)
where the RTD mechanism provides robust token-level understanding. Expanding dimensions for weaker models like SimCSE improves compatibility during ensemble integration:
$X_{wide}=\text{ExpandDims}(X_{electra}).$ (6)
III-E2 Wide Output Configurations
For models with lower baseline performance, we expanded the output dimensions:
$X_{out}^{\text{wide}}=\text{Concat}(X_{transformer},X_{context}),$ (7)
where contextual information is explicitly integrated, enhancing representation diversity.
III-E3 Bi-LSTM and Sector Contexts
Bi-LSTM layers were adapted to integrate grouped sector-level contexts:
$X_{sector}=\text{Bi-LSTM}(X_{context[0]}),$ (8)
where $context[0]$ represents sector-level information (e.g., F21 for ”F”). This hierarchical approach adds a structured representation for weakly supervised data.
III-F Loss Function
The primary loss function is the Pearson correlation loss:
$L_{pearson}=-\frac{\text{Cov}(Y_{pred},Y_{true})}{\sigma(Y_{pred})\cdot\sigma(Y_{true})},$ (9)
where Cov represents covariance, and $\sigma$ denotes standard deviation. Additionally, we employed a mean squared error (MSE) loss as a secondary measure:
$L_{mse}=\frac{1}{n}\sum_{i=1}^{n}(Y_{pred,i}-Y_{true,i})^{2}.$ (10)
These loss functions, combined with AWP and dynamic target shuffling, ensure robust optimization.
III-G Data Preprocessing
Effective data preprocessing is critical to model performance. The following steps were implemented:
III-G1 Target Grouping and Stratification
Data was grouped by anchor phrases and stratified based on semantic similarity scores:
$G=\text{GroupBy}(\text{Anchor},\text{Context})[\text{Target}],$ (11)
ensuring balanced data distribution across training folds. Targets sharing common words with anchors were allocated to the same folds to maintain contextual consistency. The Semantic Similarity Average Score Graphic and the Target Distribution Bar graph in Fig 4 show the Average Semantic Similarity Score of each combination of anchor and context, simulating the results of the grouping and stratification of the target.
Figure 4: Semantic similarity average score graph.
III-G2 Dynamic Target Shuffling
During each training step, target sequences were shuffled dynamically:
$S_{targets}^{(i)}=\text{Shuffle}(T^{(i)}),$ (12)
where $S_{targets}^{(i)}$ represents the shuffled target set at step $i$. This reduces overfitting and exposes the model to diverse input combinations.
III-G3 Contextual Augmentation
Sector-level contexts were extracted and added to the input:
$X_{input}^{aug}=\text{Concat}(X_{anchor},X_{target},X_{sector}).$ (13)
This augmentation enriches the input representation, aligning it with hierarchical domain knowledge.
III-G4 Tokenization and Padding
Inputs were tokenized using a subword tokenizer and padded to a uniform sequence length:
$X_{token}=\text{Pad}(\text{Tokenizer}(X_{raw})).$ (14)
Padding ensured compatibility with batch processing while preserving contextual integrity.
IV Evaluation Metrics
The performance of the models was evaluated using the following metrics:
IV-1 Pearson Correlation Coefficient
The primary metric is the Pearson correlation coefficient, which measures the linear correlation between predicted and true scores:
$\rho=\frac{\text{Cov}(Y_{pred},Y_{true})}{\sigma(Y_{pred})\cdot\sigma(Y_{true})},$ (15)
where Cov represents covariance, and $\sigma$ is the standard deviation.
IV-2 Mean Squared Error (MSE)
To evaluate prediction accuracy, the mean squared error was computed:
$\text{MSE}=\frac{1}{n}\sum_{i=1}^{n}(Y_{pred,i}-Y_{true,i})^{2}.$ (16)
This metric captures the average squared difference between predictions and actual values.
IV-3 F1-Score
F1-score was used to evaluate the balance between precision and recall for binary classification tasks:
[…]
This ensures a comprehensive evaluation of the model’s performance on edge cases.
IV-4 Area Under Curve (AUC)
The AUC metric evaluates the ability of the model to distinguish between classes by calculating the area under the Receiver Operating Characteristic (ROC) curve:
$\text{AUC}=\int_{0}^{1}TPR(FPR)d(FPR),$ (18)
where TPR is the true positive rate and FPR is the false positive rate.
V Experiment Results
Table I provides a detailed comparison of the proposed models and ablation studies across the evaluation metrics. The changes in model training indicators are shown in Figure 5.
Figure 5: Model indicator change chart.
TABLE I: Performance Comparison and Ablation Studies
Model Pearson (%) MSE F1-Score (%) AUC (%)
DeBERTa-v3-large 86.1 0.015 88.5 91.2
DeBERTa + LSTM 86.6 0.014 89.1 92.3
+ Linear Attention Pooling 86.8 0.013 89.4 92.8
+ Target Shuffling 87.2 0.012 90.1 93.5
Ensemble Model 87.5 0.011 91.2 94.7
VI Conclusion
This study demonstrates the effectiveness of leveraging transformer-based architectures with Bi-LSTM enhancements, adversarial weight perturbation, and dynamic preprocessing strategies for semantic similarity detection in patent documents. The integration of diverse models, combined with linear attention pooling and target shuffling, significantly improves robustness and accuracy. The ensemble strategy achieves state-of-the-art performance across multiple evaluation metrics, setting a robust foundation for practical applications in patent search and examination processes. Future work will explore domain-specific pretraining and advanced augmentation techniques to further enhance model generalization.